The function p(d)= 1+d/33, gives the pressure in atmosphere (atm), at a depth d in the sea (d is in feet). Note that p(0)= 1 atm, p (33= 2 and so on. Find the pressure at 160 feet. Write answer as a fraction.
Answer by nyc_function(2741)

p(d)= 1+d/33

Let d = 160 and solve for p(160).


p(160) = 1 + 160/33


p(160) = 1 + 4_28/33

p(160) = 5_28/33